# Beyond PCA: Manifold Dimension Estimation via Local Graph Structure

### v2

@@ 2.3 Capturing Local Graph Structure
Let $\mathcal{M}$ denote the set of all manifolds embedded in $\mathbb{R}^{p}$. A natural way to partition $\mathcal{M}$ is to group its elements according to their intrinsic dimensions. Equivalently, since any $j$-dimensional manifold can locally be expressed as the graph of smooth functions with $j$-dimensional inputs and (pj)(p-j)-dimensional outputs—whose class is denoted by $C_{j}^{\infty}$—the set $\mathcal{M}$ can be written as the disjoint union
$\mathcal{M}=\bigsqcup_{j=1}^{p}\mathcal{M}_{j}=\bigsqcup_{j=1}^{p}\left\{M\in\mathcal{M}:M\text{ is locally the graph of }C_{j}^{\infty}\text{ functions}\right\}.$
Given a dataset in $\mathbb{R}^{p}$ that satisfies the manifold hypothesis, the underlying manifold must belong to one of these subsets $\mathcal{M}_{j}$—specifically, to $\mathcal{M}_{d}$ if its intrinsic dimension is $d$. Our approach to manifold dimension estimation aims to recover the correct local graph structure of the underlying manifold across its neighborhoods. In other words, we seek to identify the function class $C_{j}^{\infty}$ from which the local representation functions originate, which best captures the relationships among the components of observed data points.
To operationalize this idea, we specify for each candidate dimension $j$ a model family $G_{j}\subseteq C_{j}^{\infty}$. The data are then fitted to the models in $G_{j}$, and a corresponding goodness-of-fit score $\sigma_{j}$ is computed to quantify how well each model family explains the data. Ideally, the model is correctly specified (or best approximated) when $j=d$, resulting in the highest score. Hence, the intrinsic dimension is estimated as
@@
In its most general form, a full model specification may span all $p$ possible partition classes. Within each partition, one may consider multiple families of $C_{j}^{\infty}$ functions, each tailored to model specific local regions of the manifold. The theory guarantees the existence of local graph representation under very mild conditions: it suffices for $g\in C_{j}^{\infty}$ to be a smooth function satisfying $\nabla g(\bm{0})=\bm{0}$. However, this level of generality is typically too broad to support practical estimation or meaningful model comparison.
To make progress, we must constrain the function classes used to fit data within each partition. The choice of these classes should ideally be guided by our prior knowledge of the underlying manifold. In the absence of such information, a natural default is to take each $G_{j}$ as the family of quadratic functions. Specifically, we approximate the output of the graph function $g\colon\mathbb{R}^{j}\to\mathbb{R}^{p-j}$ using a tensor of quadratic forms. Let $\mathcal{Q}\in\mathbb{R}^{(p-j)\times j\times j}$ be a third-order tensor, where each slice $\mathcal{Q}[\ell,:,:]=Q_{\ell}$ is a symmetric matrix encoding the second-order variation of the manifold in the $\ell$-th output direction. Then the functions in $G_{j}$ take the form
@@ 3.1 Regression with PCA Coordinates
Following the above discussion, given a data point $\bm{x}_{k}$ and its $K$-nearest neighbors $\bm{x}_{k}^{1},\ldots,\bm{x}_{k}^{K}$ (ranked by their Euclidean distances from $\bm{x}_{k}$), we first apply PCA to compute the eigendecomposition of the sample covariance matrix of these points. This yields an orthonormal basis $\bm{v}_{1},\ldots,\bm{v}_{p}$ with corresponding eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$. Taking the local mean $\bar{\bm{x}}_{k}$ as the new origin, each point in the neighborhood can then be expressed in a new coordinate system aligned with the principal directions, denoted by $\bm{x}_{0}^{\prime},\ldots,\bm{x}_{K}^{\prime}$.
@@
With the PCA coordinates, recall the quadratic models
@@
which, when cross terms among the $x_{j}$’s are included as predictors, can be expressed as a linear model. The model can then be fitted by treating the last pjp-j PCA coordinates as response variables and applying ordinary least squares (OLS).

### v3

@@ 2.3 Capturing Local Graph Structure
For $j=1,\ldots,p-1$, let $C_{j}^{\infty}$ denote the class of smooth functions with $j$-dimensional inputs and (pj)(p-j)-dimensional. Let $\mathcal{M}$ denote the set of all manifolds embedded in $\mathbb{R}^{p}$. A natural way to partition $\mathcal{M}$ is to group its elements according to their intrinsic dimensions. Equivalently, since any $j$-dimensional manifold can locally be expressed as the graph of a function in $C_{j}^{\infty}$, the set $\mathcal{M}$ can be written as the disjoint union
$\mathcal{M}=\bigsqcup_{j=1}^{p}\mathcal{M}_{j},$
where $\mathcal{M}_{j}$ is the set of all manifolds $M\in\mathcal{M}_{j}$ that can be locally represented around every point $\bm{x}_{0}\in M$ as the graph of a map $g_{j,\bm{x}_{0}}\in C_{j}^{\infty}$, as in equation (1). Collecting all such local representations gives
$\mathcal{G}_{j}:=\bigcup_{M\in\mathcal{M}_{j}}\left\{g_{j,\bm{x}_{0}}:\bm{x}_{0}\in M\right\}\subseteq C_{j}^{\infty},$
which can be viewed as the space of all smooth local graph representations of $j$-dimensional manifolds in $\mathbb{R}^{p}$.
Given a dataset in $\mathbb{R}^{p}$ that satisfies the manifold hypothesis, its unknown underlying manifold must belong to one of these subsets $\mathcal{M}_{j}$—specifically, to $\mathcal{M}_{d}$ if its intrinsic dimension is dd—and thus be locally represented by $\mathcal{G}_{d}$. Our approach to manifold dimension estimation is to identify $\mathcal{G}_{d}$.
To operationalize this idea, we specify for each candidate dimension $j$ a family $G_{j}\subseteq C_{j}^{\infty}$ of (parametric) models assumed to be close enough to the functions in $\mathcal{G}_{j}$. The data are then fitted to the models in $G_{j}$, and a corresponding goodness-of-fit score $\sigma_{j}$ is computed to quantify how well each model family explains the data. Ideally, the model is correctly specified (or best approximated) when $j=d$, resulting in the highest score. Hence, the intrinsic dimension is estimated as
@@
In its most general form, a full model specification may span all $p$ possible partition classes. Within each partition, one may consider multiple families of functions from $C_{j}^{\infty}$, each tailored to model specific local regions of the manifold. The theory guarantees the existence of local graph representation under very mild conditions: it suffices for $g\in C_{j}^{\infty}$ to be a smooth function satisfying $\nabla g(\bm{0})=\bm{0}$. However, this level of generality is typically too broad to support practical estimation or meaningful model comparison.
To make progress, we must constrain the function classes used to fit data within each partition. The choice of these classes should ideally be guided by our prior knowledge of the underlying manifold. In the absence of such information, a natural default is to take each $G_{j}$ as a family of quadratic functions. Specifically, we approximate the output of the graph function $g\colon\mathbb{R}^{j}\to\mathbb{R}^{p-j}$ using a tensor of quadratic forms. Let $\mathcal{Q}\in\mathbb{R}^{(p-j)\times j\times j}$ be a third-order tensor, where each slice $\mathcal{Q}[\ell,:,:]=Q_{\ell}$ is a symmetric matrix encoding the second-order variation of the manifold in the $\ell$-th output direction. Then the functions in $G_{j}$ take the form
@@ 3.1 Regression with PCA Coordinates
Following the above discussion, given a data point $\bm{x}_{k}\in\{\bm{x}_{1},\ldots,\bm{x}_{n}\}$ and its $K$-nearest neighbors $\bm{x}_{k}^{1},\ldots,\bm{x}_{k}^{K}$ (ranked by their Euclidean distances from $\bm{x}_{k}$), we first apply PCA to compute the eigendecomposition of the sample covariance matrix of these $K+1$ points (after centering them at their local mean, $\bar{\bm{x}}_{k})$. This yields an orthonormal basis whose vectors $\bm{v}_{1},\ldots,\bm{v}_{p}$ are stored in matrix $V$, with corresponding eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$. Taking the local mean $\bar{\bm{x}}_{k}$ as the new origin, each point in the neighborhood can then be expressed in a new coordinate system aligned with the principal directions, denoted by $\bm{x}_{0}^{\prime},\ldots,\bm{x}_{K}^{\prime}$, which are given by
$\bm{x}_{0}^{\prime}=V^{\top}(\bm{x}_{k}-\bar{\bm{x}}_{k}),\quad\bm{x}_{\ell}^{\prime}=V^{\top}(\bm{x}_{k}^{\ell}-\bar{\bm{x}}_{k}),\ \ \ \ell=1,\cdots,K.$
@@
Using the PCA coordinates, recall the quadratic models
@@
which can be rewritten as linear models by expanding all quadratic and cross-product terms of the inputs $\bm{x}^{\prime}_{1},\ldots,\bm{x}^{\prime}_{j}$. For each neighborhood, multivariate linear models such as
$\bm{x}_{(j+1):p}^{\prime}=(q_{1}(\bm{x}_{1:j}^{\prime}),\ldots,q_{p-j}(\bm{x}_{1:j}^{\prime}))^{\top}+\bm{\epsilon}$
($j=1,\ldots,p-1$) could then be fitted to recover (1) by treating the last pjp-j PCA coordinates as response variables and applying ordinary least squares (OLS).
Our QE estimator simplifies this approach by restricting ourselves to univariate linear models
$x_{j+1}^{\prime}=q_{1}(\bm{x}_{1:j}^{\prime})+\epsilon.$
This has two key advantages: first, we avoid PCA coordinates associated with smaller eigenvalues which are typically more sensitive to noise; second, it helps us avoid spurious relationships between input and output when $j<d$. In contrast, using all of the last pjp-j coordinates as responses introduces the risk of detecting artificial dependencies, which can negatively affect model evaluation.